# Central Limit Theorems for Asynchronous Averaged Q-Learning

### v1

@@ Abstract
This paper establishes central limit theorems for Polyak–Ruppert averaged Q-learning under asynchronous updates. We present a non-asymptotic central limit theorem, where the convergence rate in Wasserstein distance explicitly reflects the dependence on the number of iterations, state–action space size, the discount factor, and the quality of exploration. In addition, we derive a functional central limit theorem, showing that the partial-sum process converges weakly to a Brownian motion.
@@ 1 Introduction
Reinforcement Learning (RL) has emerged as a powerful paradigm in artificial intelligence, achieving successes in various applications such as Atari [Mnih et al.(2015)Mnih, Kavukcuoglu, Silver, Rusu, Veness, Bellemare, Graves, Riedmiller, Fidjeland, Ostrovski, et al.], Go [Silver et al.(2016)Silver, Huang, Maddison, Guez, Sifre, Van Den Driessche, Schrittwieser, Antonoglou, Panneershelvam, Lanctot, et al.], robot manipulation [Tan et al.(2018)Tan, Zhang, Coumans, Iscen, Bai, Hafner, Bohez, and Vanhoucke, Zeng et al.(2020)Zeng, Song, Lee, Rodriguez, and Funkhouser], and aligning large language models to human preferences [Shao et al.(2024)Shao, Wang, Zhu, Xu, Song, Bi, Zhang, Zhang, Li, Wu, et al., Ouyang et al.(2022)Ouyang, Wu, Jiang, Almeida, Wainwright, Mishkin, Zhang, Agarwal, Slama, Ray, et al.]. Q-learning [Watkins and Dayan(1992)], which directly learns the optimal action-value function (Q-function) from experience trajectories, is one of the most widely used RL algorithms.
@@
Stochastic approximation (SA) [Benveniste et al.(2012)Benveniste, Métivier, and Priouret, Borkar and Borkar(2008)] is a general iterative framework to solve fixed-point equation problems. Since the Bellman operator in RL is a contraction map with a unique fixed point, many RL algorithms can be interpreted as instances of SA. For example, TD learning [Sutton and Barto(1998)] can be viewed as an instance of linear SA. Synchronous Q-learning, by contrast, is a special case of nonlinear SA with martingale noise. The asynchronous Q-learning algorithm studied in this work, however, is a nonlinear SA problem with Markovian noise. There is a growing line of work on finite-sample analysis of SA with applications to RL algorithms [Wainwright(2019), Srikant and Ying(2019), Bhandari et al.(2018)Bhandari, Russo, and Singal, Chen et al.(2022)Chen, Zhang, Doan, Clarke, and Maguluri, Chen et al.(2020b)Chen, Maguluri, Shakkottai, and Shanmugam, Qu and Wierman(2020), Khodadadian and Zubeldia(2025), Chandak et al.(2025)Chandak, Haque, and Bambos, Chandak(2025)].
@@
Polyak-Ruppert averaging is a classical variance-reduction technique to stabilize and accelerate SA algorithms. In this paper, we are interested in establishing central limit theorems (CLTs) for Polyak-Ruppert averaged Q-learning under asynchronous updates. Building CLTs provides a foundational understanding of the algorithm’s statistical properties. This asymptotic normality is crucial for uncertainty quantification and statistical inference in RL. Building on the seminal work by Polyak and Juditsky [Polyak and Juditsky(1992)], a non-asymptotic CLT for Polyak-Ruppert averaged SGD was established [Anastasiou et al.(2019)Anastasiou, Balasubramanian, and Erdogdu]. [Mou et al.(2020)Mou, Li, Wainwright, Bartlett, and Jordan, Samsonov et al.(2025)Samsonov, Sheshukova, Moulines, and Naumov] derive non-asymptotic CLTs for linear SA with Polyak–Ruppert averaged iterates. Similar results for two-time-scale SA are also studied [Han et al.(2024)Han, Li, Liang, and Zhang, Hu et al.(2024)Hu, Doshi, et al., Kong et al.(2025)Kong, Zeng, Doan, and Srikant]. Recently, CLTs for SA with applications to RL algorithms are studied [Borkar et al.(2025)Borkar, Chen, Devraj, Kontoyiannis, and Meyn]. As a special case linear SA, [Srikant(2024), Samsonov et al.(2024)Samsonov, Moulines, Shao, Zhang, and Naumov] derive non-asymptotic CLTs for TD-learning with averaging. However, non-asymptotic CLTs for Q-learning remain unexplored.
@@
As a special case of nonlinear SA, Q-learning is substantially more challenging to analyze than linear SA and TD learning. Functional CLTs for Polyak–Ruppert averaged synchronous Q-learning was established in [Xie and Zhang(2022), Li et al.(2023)Li, Yang, Liang, Zhang, and Jordan, Panda et al.(2025)Panda, Liu, and Xiang]. Synchronous Q-learning only considers martingale noises. By contrast, asynchronous Q-learning updates a single state–action pair based on one transition sample at each iteration, which involves Markovian noises that are non-IID. Moreover, the empirical Bellman operator in synchronous Q-learning is smooth, whereas in asynchronous Q-learning, it is non-smooth. Thus, the challenges in analyzing asynchronous Q-learning come from nonlinearity, Markovian samples, and a non-smooth operator. Recently, [Zhang and Xie(2024)] established a functional CLT for asynchronous Q-learning with a constant stepsize. Constant stepsize does not satisfy the necessary conditions for establishing a non-asymptotic CLT, which we detail in Section 3. To the best of our knowledge, no non-asymptotic CLT is currently known for Q-learning, even in the synchronous setting. In this work, we close this gap and prove both a non-asymptotic CLT and a functional CLT for asynchronous averaged Q-learning with decaying stepsizes.
@@ 2 Preliminaries
$\Gamma$ is the temporal difference in the Q-function iterate. The sample trajectory $\{(s_{k},a_{k})\}$ is collected by the MDP under a behavior policy $\pi_{b}$. We define $V_{k}(s):=\max_{a}Q_{k}(s,a)$. Now we make the following assumption on the Markov chain, which is standard in the literature [Zhang and Xie(2024), Zhang et al.(2022)Zhang, Des Combes, and Laroche, Chen et al.(2021)Chen, Maguluri, Shakkottai, and Shanmugam, Li et al.(2020)Li, Wei, Chi, Gu, and Chen, Qu and Wierman(2020)].
@@
Lemma 2.1 (Proposition 3.1 in [Chen et al.(2021)Chen, Maguluri, Shakkottai, and Shanmugam]).
@@ 3.1 Non-Asymptotic Central Limit Theorem
The stepsize in the Q-learning update chosen in this work is $\alpha_{k}=\alpha(k+b)^{-\beta}$ for two reasons. First, convergence of stochastic approximation with averaging schemes relies on several key conditions [Polyak and Juditsky(1992), Li et al.(2023)Li, Yang, Liang, Zhang, and Jordan]: (i) $0\leq\sup_{k}\alpha_{k}\leq 1$, $\alpha_{k}\downarrow 0$and $k\alpha_{k}\uparrow\infty$; (ii) $\frac{\alpha_{k-1}-\alpha_{k}}{\alpha_{k-1}}=o(\alpha_{k-1})$; (iii) $\frac{1}{\sqrt{K}}\sum_{k=0}^{K}\alpha_{k}\rightarrow 0$; (iv) $\frac{\sum_{k=0}^{K}\alpha_{k}}{K\alpha_{K}}\leq C$. Constant stepsizes violate conditions (i) and (iii), while linear stepsizes violate condition (ii). By contrast, polynomial stepsizes satisfy all of the above. Second, the problem-dependent constants $\alpha$ and $b$ are crucial for establishing a finite-sample convergence guarantee [Chen et al.(2021)Chen, Maguluri, Shakkottai, and Shanmugam], which we leverage in our analysis. The parameter $\alpha$ acts as a scaling factor for balancing the trade-off between the speed of convergence and the final error of the algorithm. The parameter $b$ is used to control the magnitude of the initial stepsizes and ensure the stability of the algorithm during the early stages. Setting $\beta=2/3$, the rate of convergence can be simplified as follows.
@@ 3.2 Functional Central Limit Theorem
The FCLT is an important extension to the conventional CLT. Donsker’s FCLT [Donsker(1951)] states that the normalized partial sum process of i.i.d. random variables converges weakly to a Brownian motion in the Skorokhod space. In this section, we establish an FCLT for asynchronous Q-learning iterates, showing that the partial-sum process converges in distribution to a rescaled Brownian motion. Let $\mathcal{D}[0,1]$ denote the Skorokhod space. For $\zeta\in[0,1]$, we define the standardized partial sum processes associated with $\{Q_{k}\}_{k\geq 1}$ as
@@
We can see that the conventional CLT is a special case of the FCLT when $\zeta=1$. As the FCLT provides a basis for the asymptotic normality of certain functionals of stochastic processes, it is important for uncertainty quantification and statistical inference in Q-learning. Previous works have established the FCLT for synchronous Q-learning [Xie and Zhang(2022), Li et al.(2023)Li, Yang, Liang, Zhang, and Jordan, Panda et al.(2025)Panda, Liu, and Xiang]. A recent work [Zhang and Xie(2024)] established a FCLT for asynchronous Q-learning with a constant step size. In contrast, our result concerns diminishing step-sizes.
@@ 4 Conclusion
Next, we prove $\Delta^{\downarrow}_{k}\leq\Delta_{k}\leq\Delta^{\uparrow}_{k}$ by induction. The base case $k=0$ holds by initialization. Suppose the statement holds at $k$. We observe that, since $\alpha_{k}$ and the entries in matrices $D$ and $P^{\pi^{*}}$ are all bounded between 0 and $1$, the entries in matrix $I-\alpha_{k}D+\alpha_{k}\gamma DP^{\pi^{*}}$ are nonnegative. Consequently,
@@
We decompose this term into two parts, where the first part has a bounded norm and the second part is a bounded martingale difference sequence. To this end, we use the Poisson equation technique [Glynn and Meyn(1996), Douc et al.(2018)Douc, Moulines, Priouret, Soulier, Douc, Moulines, Priouret, and Soulier, Makowski and Shwartz(2002), Chen et al.(2020a)Chen, Devraj, Busic, and Meyn] to transform the Markovian noise into a martingale difference sequence. By a standard use of the technique [Douc et al.(2018)Douc, Moulines, Priouret, Soulier, Douc, Moulines, Priouret, and Soulier], we know there exists a solution $X_{k}:\tilde{\mathcal{S}}\rightarrow\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}$ to the following Poisson’s equation for all $k\in[K]$,
@@
where $\phi_{6}$ is a bounded martingale difference sequence. Note we have also established bounds for $\{\phi_{i}\}_{i=1}^{5}$. Therefore, to establish CLTs for the averaged $Q$-learning iterates, we can apply any suitable known martingale CLTs. To proceed, we choose the non-asymptotic martingale CLT given in [Srikant(2024)]. We prove in Lemma A.5 that $\mathcal{W}_{1}\left(\frac{1}{\sqrt{K}}\phi_{6},(A^{-1}\Sigma A^{-\top})^{1/2}\mathcal{N}(0,I))\right)\leq O\left(\frac{1}{(1-\gamma)\rho K^{\beta/2}}\right)$. Note that
@@ A.3 Martingale CLT
Theorem A.4 (Restatement of Theorem 1 in [Srikant(2024)]).
Let $\{m_{k}\}_{k\geq 1}$ be a $d$-dimensional martingale difference sequence with respect to a filtration $\{\mathcal{F}_{k}\}_{k\geq 0}$. Assume (i) $\mathbb{E}[\|m_{k}\|_{2}]\leq\infty$ and $\mathbb{E}[m_{k}|\mathcal{F}_{k-1}]=0$ for all $k\geq 1$; (ii) $\mathbb{E}[\|m_{k}\|_{2}^{2+\beta}]$ exists almost surely for all $k\geq 1$ and some $\beta\in(0,1)$ and (iii) $\Sigma_{\infty}=\mathbb{E}[m_{k}m_{k}^{\top}|\mathcal{F}_{k-1}]$ exists and further assume that $\lim_{n\rightarrow\infty}(\Sigma_{1}+\cdots+\Sigma_{n})/n=\Sigma_{\infty}$ almost surely for some positive definite $\Sigma_{\infty}$. It follows that
@@
$\tilde{\Sigma}_{k}=\sum_{i,j\in\tilde{S}}\tilde{p}_{i}(k)\tilde{P}(i,j)(X_{k}(j)-\mathbb{E}[X_{k}(Y_{1})|Y_{0}=i])(X_{k}(j)-\mathbb{E}[X_{k}(Y_{1})|Y_{0}=i])^{\top}$
@@
Now we can substitute $n=K$, $m_{k}=A^{-1}(X_{k-1}(Y_{k})-\mathbb{E}[X_{k-1}(Y_{k})|Y_{k-1}])$, $\Sigma_{k}=A^{-1}\tilde{\Sigma}_{k}A^{-\top}$, and $\Sigma_{\infty}=A^{-1}\Sigma A^{-\top}$ into Theorem A.4. Note that under Assumption 2, the three conditions in Theorem A.4 are satisfied. The rest of the proof follows from the proof of Theorem 2 in [Srikant(2024)], with necessary modifications to accommodate our setting. To conclude, with the substitutions, we have
@@
Polish space is a separable and complete function space. It is a crucial structure for applying convergence in distribution results such as FCLT. Recall that we denote $\mathcal{D}[0,1]$ as the Skorokhod space. Equipped with the Skorokhod $J_{1}$ topology with a particular metric [Prokhorov(1956)], $\mathcal{D}[0,1]$ is a Polish space. We use $\overset{\mathrm{w}}{\rightarrow}$ to denote weak convergence for some sequence of random elements. To prove the theorem, we need the following result.
@@
From the proof of Theorem 3.1, we know $\sup_{\zeta\in[0,1]}\left\|\frac{1}{\sqrt{K}}\phi_{i}(\zeta)\right\|_{\infty}=o(1)$ for $i\in\{1,2,3,4,5\}$. Let $X$ and $\Sigma$ as defined in the proof of Lemma A.5. The following lemma, which establishes the FCLT for $\frac{1}{\sqrt{K}}\phi_{6}(\zeta)$, is a direct consequence of Theorem 4.2 in [Hall and Heyde(2014)].

### v2

@@ Abstract
This paper establishes central limit theorems for Polyak–Ruppert averaged Q-learning under asynchronous updates. We prove a non-asymptotic central limit theorem, where the convergence rate in Wasserstein distance explicitly reflects the dependence on the number of iterations, state–action space size, the discount factor, and the quality of exploration. In addition, we derive a functional central limit theorem, showing that the partial-sum process converges weakly to a Brownian motion.
@@ 1 Introduction
Reinforcement Learning (RL) has emerged as a powerful paradigm in artificial intelligence, achieving successes in various applications such as Atari (Mnih et al., 2015), Go (Silver et al., 2016), robot manipulation (Tan et al., 2018; Zeng et al., 2020), and aligning large language models to human preferences (Shao et al., 2024; Ouyang et al., 2022). Q-learning (Watkins and Dayan, 1992), which directly learns the optimal action-value function (Q-function) from experience trajectories, is one of the most widely used RL algorithms.
@@
Stochastic approximation (SA) (Benveniste et al., 2012; Borkar and Borkar, 2008) is a general iterative framework to solve fixed-point equation problems. Since the Bellman operator in RL is a contraction map with a unique fixed point, many RL algorithms can be interpreted as instances of SA. For example, TD learning (Sutton and Barto, 1998) can be viewed as an instance of linear SA. Synchronous Q-learning, by contrast, is a special case of nonlinear SA with martingale noise. The asynchronous Q-learning algorithm studied in this work, however, is a nonlinear SA problem with Markovian noise. There is a growing line of work on finite-sample analysis of SA with applications to RL algorithms (Wainwright, 2019; Srikant and Ying, 2019; Bhandari et al., 2018; Chen et al., 2022, 2020b; Qu and Wierman, 2020; Khodadadian and Zubeldia, 2025; Chandak et al., 2025; Chandak, 2025).
@@
Polyak-Ruppert averaging is a classical variance-reduction technique to stabilize and accelerate SA algorithms. In this paper, we are interested in establishing central limit theorems (CLTs) for Polyak-Ruppert averaged Q-learning under asynchronous updates. Building CLTs provides a foundational understanding of the algorithm’s statistical properties. This asymptotic normality is crucial for uncertainty quantification and statistical inference in RL. Building on the seminal work by Polyak and Juditsky (1992), a non-asymptotic CLT for Polyak-Ruppert averaged SGD was established (Anastasiou et al., 2019). Mou et al. (2020); Samsonov et al. (2025) derive non-asymptotic CLTs for linear SA with Polyak–Ruppert averaged iterates. Similar results for two-time-scale SA are also studied (Han et al., 2024; Hu et al., 2024; Kong et al., 2025). Recently, CLTs for SA with applications to RL algorithms are studied (Borkar et al., 2025). As a special case linear SA, Srikant (2024); Samsonov et al. (2024) derive non-asymptotic CLTs for TD-learning with averaging. However, non-asymptotic CLTs for Q-learning remain unexplored.
@@
As a special case of nonlinear SA, Q-learning is substantially more challenging to analyze than linear SA and TD learning. Functional CLTs for Polyak–Ruppert averaged synchronous Q-learning was established in Xie and Zhang (2022); Li et al. (2023); Panda et al. (2025). Synchronous Q-learning only considers martingale noises. By contrast, asynchronous Q-learning updates a single state–action pair based on one transition sample at each iteration, which involves Markovian noises that are non-IID. Moreover, the empirical Bellman operator in synchronous Q-learning is smooth, whereas in asynchronous Q-learning, it is non-smooth. Thus, the challenges in analyzing asynchronous Q-learning come from nonlinearity, Markovian samples, and a non-smooth operator. Recently, Zhang and Xie (2024) established a functional CLT for asynchronous Q-learning with a constant stepsize. Constant stepsize does not satisfy the necessary conditions for establishing a non-asymptotic CLT, which we detail in Section 3. To the best of our knowledge, no non-asymptotic CLT is currently known for Q-learning, even in the synchronous setting. In this work, we close this gap and prove both a non-asymptotic CLT and a functional CLT for asynchronous averaged Q-learning with decaying stepsizes.
@@ 2 Preliminaries
$\Gamma$ is the temporal difference in the Q-function iterate. The sample trajectory $\{(s_{k},a_{k})\}$ is collected by the MDP under a behavior policy $\pi_{b}$. We define $V_{k}(s):=\max_{a}Q_{k}(s,a)$. Now we make the following assumption on the Markov chain, which is standard in the literature (Zhang and Xie, 2024; Zhang et al., 2022; Chen et al., 2021; Li et al., 2020; Qu and Wierman, 2020).
@@
Lemma 2.1 (Proposition 3.1 in Chen et al. (2021)).
@@ 3.1 Non-Asymptotic Central Limit Theorem
The stepsize in the Q-learning update chosen in this work is $\alpha_{k}=\alpha(k+b)^{-\beta}$ for two reasons. First, convergence of stochastic approximation with averaging schemes relies on several key conditions (Polyak and Juditsky, 1992; Li et al., 2023): (i) $0\leq\sup_{k}\alpha_{k}\leq 1$, $\alpha_{k}\downarrow 0$ and $k\alpha_{k}\uparrow\infty$; (ii) $\frac{\alpha_{k-1}-\alpha_{k}}{\alpha_{k-1}}=o(\alpha_{k-1})$; (iii) $\frac{1}{\sqrt{K}}\sum_{k=0}^{K}\alpha_{k}\rightarrow 0$; (iv) $\frac{\sum_{k=0}^{K}\alpha_{k}}{K\alpha_{K}}\leq C$. Constant stepsizes violate conditions (i) and (iii), while linear stepsizes violate condition (ii). By contrast, polynomial stepsizes satisfy all of the above. Second, the problem-dependent constants $\alpha$ and $b$ are crucial for establishing a finite-sample convergence guarantee (Chen et al., 2021), which we leverage in our analysis. The parameter $\alpha$ acts as a scaling factor for balancing the trade-off between the speed of convergence and the final error of the algorithm. The parameter $b$ is used to control the magnitude of the initial stepsizes and ensure the stability of the algorithm during the early stages. Setting $\beta=2/3$, the rate of convergence can be simplified as follows.
@@ 3.2 Functional Central Limit Theorem
The FCLT is an important extension to the conventional CLT. Donsker’s FCLT (Donsker, 1951) states that the normalized partial sum process of i.i.d. random variables converges weakly to a Brownian motion in the Skorokhod space. In this section, we establish an FCLT for asynchronous Q-learning iterates, showing that the partial-sum process converges in distribution to a rescaled Brownian motion. Let $\mathcal{D}[0,1]$ denote the Skorokhod space. For $\zeta\in[0,1]$, we define the standardized partial sum processes associated with $\{Q_{k}\}_{k\geq 1}$ as
@@
We can see that the conventional CLT is a special case of the FCLT when $\zeta=1$. As the FCLT provides a basis for the asymptotic normality of certain functionals of stochastic processes, it is important for uncertainty quantification and statistical inference in Q-learning. Previous works have established the FCLT for synchronous Q-learning (Xie and Zhang, 2022; Li et al., 2023; Panda et al., 2025). A recent work (Zhang and Xie, 2024) established a FCLT for asynchronous Q-learning with a constant step size. In contrast, our result concerns diminishing step-sizes.
@@ 4 Conclusion
Next, we prove $\Delta^{\downarrow}_{k}\leq\Delta_{k}\leq\Delta^{\uparrow}_{k}$ by induction. The base case $k=0$ holds by initialization. Suppose the statement holds at $k$. We observe that, since $\alpha_{k}$ and the entries in matrices $D$ and $P^{\pi^{*}}$ are all bounded between 0 and 11, the entries in matrix $I-\alpha_{k}D+\alpha_{k}\gamma DP^{\pi^{*}}$ are nonnegative. Consequently,
@@
We decompose this term into two parts, where the first part has a bounded norm and the second part is a bounded martingale difference sequence. To this end, we use the Poisson equation technique (Glynn and Meyn, 1996; Douc et al., 2018; Makowski and Shwartz, 2002; Chen et al., 2020a) to transform the Markovian noise into a martingale difference sequence. By a standard use of the technique (Douc et al., 2018), we know there exists a solution $X_{k}:\tilde{\mathcal{S}}\rightarrow\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}$ to the following Poisson’s equation for all $k\in[K]$,
@@
where $\phi_{6}$ is a bounded martingale difference sequence. Note we have also established bounds for $\{\phi_{i}\}_{i=1}^{5}$. Therefore, to establish CLTs for the averaged $Q$-learning iterates, we can apply any suitable known martingale CLTs. To proceed, we choose the non-asymptotic martingale CLT given in Srikant (2024). We prove in Lemma A.5 that $\mathcal{W}_{1}\left(\frac{1}{\sqrt{K}}\phi_{6},(A^{-1}\Sigma A^{-\top})^{1/2}\mathcal{N}(0,I))\right)\leq O\left(\frac{1}{(1-\gamma)\rho K^{\beta/2}}\right)$. Note that
@@ A.3 Martingale CLT
Theorem A.4 (Restatement of Theorem 1 in Srikant (2024)).
Let $\{m_{k}\}_{k\geq 1}$ be a $d$-dimensional martingale difference sequence with respect to a filtration $\{\mathcal{F}_{k}\}_{k\geq 0}$. Assume (i) $\mathbb{E}[\|m_{k}\|_{2}]\leq\infty$ and $\mathbb{E}[m_{k}|\mathcal{F}_{k-1}]=0$ for all $k\geq 1$; (ii) $\mathbb{E}[\|m_{k}\|_{2}^{2+\beta}]$ exists almost surely for all $k\geq 1$ and some $\beta\in(0,1)$ and (iii) $\Sigma_{k}=\mathbb{E}[m_{k}m_{k}^{\top}|\mathcal{F}_{k-1}]$ exists and further assume that $\lim_{n\rightarrow\infty}(\Sigma_{1}+\cdots+\Sigma_{n})/n=\Sigma_{\infty}$ almost surely for some positive definite $\Sigma_{\infty}$. It follows that
@@
$\tilde{\Sigma}_{k}=\sum_{i,j\in\tilde{S}}\tilde{p}_{k}(i)\tilde{P}(i,j)(X_{k}(j)-\mathbb{E}[X_{k}(Y_{1})|Y_{0}=i])(X_{k}(j)-\mathbb{E}[X_{k}(Y_{1})|Y_{0}=i])^{\top}$
@@
Now we can substitute $n=K$, $m_{k}=A^{-1}(X_{k-1}(Y_{k})-\mathbb{E}[X_{k-1}(Y_{k})|Y_{k-1}])$, $\Sigma_{k}=A^{-1}\tilde{\Sigma}_{k}A^{-\top}$, and $\Sigma_{\infty}=A^{-1}\Sigma A^{-\top}$ into Theorem A.4. Note that under Assumption 2, the three conditions in Theorem A.4 are satisfied. The rest of the proof follows from the proof of Theorem 2 in Srikant (2024), with necessary modifications to accommodate our setting. To conclude, with the substitutions, we have
@@
Polish space is a separable and complete function space. It is a crucial structure for applying convergence in distribution results such as FCLT. Recall that we denote $\mathcal{D}[0,1]$ as the Skorokhod space. Equipped with the Skorokhod $J_{1}$ topology with a particular metric (Prokhorov, 1956), $\mathcal{D}[0,1]$ is a Polish space. We use $\overset{\mathrm{w}}{\rightarrow}$ to denote weak convergence for some sequence of random elements. To prove the theorem, we need the following result.
@@
From the proof of Theorem 3.1, we know $\sup_{\zeta\in[0,1]}\left\|\frac{1}{\sqrt{K}}\phi_{i}(\zeta)\right\|_{\infty}=o(1)$ for $i\in\{1,2,3,4,5\}$. Let $X$ and $\Sigma$ as defined in the proof of Lemma A.5. The following lemma, which establishes the FCLT for $\frac{1}{\sqrt{K}}\phi_{6}(\zeta)$, is a direct consequence of Theorem 4.2 in Hall and Heyde (2014).